OpenStudy (anonymous):

if a and b are roots of the quadratic equation ax^2+bx+c=0, then quadratic equation ax^2-bx(x-1)+c(x-1)^2=0 has roots ?

OpenStudy (anonymous):

@Abhisar ? @bek_love ? @Compassionate ? @DollyAcquah ? @emmigrace222 ? @ganeshie8 ?

OpenStudy (emmigrace222):

Well then

OpenStudy (anonymous):

how to proceed ?

OpenStudy (anonymous):

sorry the roots are

OpenStudy (anonymous):

\[\alpha,\beta\]

OpenStudy (emmigrace222):

honestly i dont know

OpenStudy (anonymous):

i wrote \[ax^2+bx+c=a(x-\alpha)(x-\beta)\]

OpenStudy (emmigrace222):

thats really good :)

OpenStudy (anonymous):

then how to proceed?

OpenStudy (anonymous):

ax^2-bx(x-1)+c(x-1)^2=0

OpenStudy (anonymous):

should i expand it?

OpenStudy (anonymous):

like it would be \[x^2(a-b+c)+x(b-2c)+c=0\]

OpenStudy (anonymous):

or\[a(x-\alpha)(x-\beta)-bx^2+cx^2-2cx=0\]

OpenStudy (anonymous):

@myininaya ? @mathmale ?

OpenStudy (anonymous):

@nincompoop ?

OpenStudy (anonymous):

@oliviagutz ?

OpenStudy (anonymous):

@Hotchellerae21 ?

OpenStudy (anonymous):

@Future_UsArmy_MP ?

OpenStudy (anonymous):

@Questionmachine ?

OpenStudy (anonymous):

i am here

OpenStudy (anonymous):

thanks can u suggest some way i have solved little bit bu how to proceed further?

OpenStudy (anonymous):

ok what do u have so far

OpenStudy (anonymous):

scroll upwards!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

maybe this site will help

OpenStudy (anonymous):

@ganeshie8 ? can u suggest also please thanks @Hotchellerae21 for ur effort

OpenStudy (anonymous):

oh ok no problem @cody_123

OpenStudy (anonymous):

@Abhisar ? @beccamarx19 ? @Compassionate ? @dan815 ? @Future_UsArmy_MP ? @Hero ?

ganeshie8 (ganeshie8):

\(ax^2+bx+c=0\) \(c/a = p \implies c = a\alpha\) \(-b/a = q \implies b = -a\beta\)

ganeshie8 (ganeshie8):

\(\large ax^2-bx(x-1)+c(x-1)^2= x^2(a-b+c) + x(b-2c) + c \)

ganeshie8 (ganeshie8):

eliminate b, c by substituting their values ^

OpenStudy (anonymous):

have you taken \[\alpha+\beta=p=-b/a\]?

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@ganeshie8 ?